Find the equation of the line with slope equal to 1 and normal to the
curve x
2 + 3xy + y
2 = 5 at P0(x0, y0) when x0 = 1.
1
Expert's answer
2021-10-19T23:49:01-0400

Solution;

Given the equation of the curve is;

"x^2+3xy+y^2=5"

At "P_0(x_0,y_0)" When "x_0=1"

Differentiate as follows;

"2x+3y+3x\frac{dy}{dx}+2y\frac{dy}{dx}=0"

Group;

"(3x+2y)\frac{dy}{dx}=-(2x+3y)"

"\frac{dy}{dx}=\frac{-(2x+3y)}{3x+2y}"

Since "x_0=1" ,by direct substitute into the equation,find values of y.

"x^2+3xy+y^2=5"

"(1^2)+3(1)y+y^2=5"

"y^2+3y-4=0"

"y^2-y+4y-4=0"

"y(y-1)+4(y-1)=0"

Hence ;

"y=1 or -4"

Use y=1 because at y=-4,the gradient of the equation is not equal to 1.

So the point "P_0(x_0,y_0)=P_0(1,1)"

"\frac{dy}{dx}(1,1)=\frac{-(2+3)}{3+2}=-1"

The equation of the normal to curve is ;

"(x-1)=1(y-1)"

"x-y=0"

This is the required solution.